A 5.0L bomb calorimeter, with a heat capacity of 14.25 kJ/°C, is filled to a pressure of 5.0 atm with O2 gas at 32.9°C. A 6.55 g sample of aniline (C6H5NH2, molar mass = 93.13 g/mol) was then combusted in this calorimeter.

Determine the value of the final temperature of the calorimeter.

4 C6H5NH2(l) + 35 O2(g) → 24 CO2(g) + 7 H2O(g) + 2 NO2(g)

ΔH°rxn= -1.28 x 104kJ/mol
Answer #1

Hi,

Hope you are doing well.

Given that,

Volume,V = 5 L = 5 x 10-3 m

Pressure, P= 5 atm = 5 x 100 Pa

Temperature, T = 32.9 + 273 K = 304.9 K

Hence, no.of moles of oxygen gas is given by,

1 PV 5 x 10° Pa x 5 x 10-3 m RT – 8.314 J/mol. K x 304.9 K

on=0.983 mol

Also, no.of moles of aniline is given by,

6.55 g ni = M - 93.13 g/mol = 0.0703 mol

Combustion reaction of aniline is given by,

4 C_{6}H_{5}NH_{2}_{(l)} + 35 O_{2}_{(g) }\rightarrow 24 CO_{2}_{(g)} + 7 H_{2}O_{(g)} + 2 NO_{2}_{(g)}

From the above balanced reaction we have, 4 moles of aniline reacts with 35 moles of O2.

Hence, 1 mole of aniline reacts with 35/4 moles of O2.

Therefore, 0.0703 moles of aniline reacts with,

35 n = 0.0703 mol x

n = 0.6154 mol

Hence, energy released for this reaction is given by,

H= 1.28 x 104 kJ/mol 35 1.28 x 104 kJ/mol *0.6154 mol 35

\mathbf{\therefore H=225.062\;kJ}

Hence, Change/increase in temperature of the calorimeter is given by,

\Delta T=\frac{H}{C}=\frac{225.062\;J}{14.25\;kJ/^{o}C}

\therefore \Delta T=15.79^{o}C}

Hence,Final temperature of the calorimeter is given by,

\therefore T_{final}=T+\Delta T=32.9^{o}C+15.79^{o}C}

· Tfinal = 48.69°C